rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_red
icon_orange
icon_green
icon_green
icon_green
icon_green
icon_blue
icon_green
icon_orange
icon_red
icon_green
icon_red
icon_red
icon_green
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#2818 — FS#6778 — bea13 cor13 dra13 fer13 gar13 jol13 men13 pra13 smg13 sml13 stm13

Attached to Project— ADSL / SDSL
Maintenance
Backend / Core
CLOSED
100%
We have a problem on the communication between the net center and the first NRA.

We are looking for the origin of the problem.
Date:  Friday, 25 May 2012, 12:44PM
Reason for closing:  Done
Comment by OVH - Thursday, 24 May 2012, 21:53PM

The tickets were opened with different operators.


Comment by OVH - Thursday, 24 May 2012, 21:56PM

The first tests show a problem on the fiber
section. The technicians will be on the spot
at 20:30 and we will have more informations
about the interruption.


Comment by OVH - Thursday, 24 May 2012, 22:44PM

The interruption was localized. The teams are on the spot.


Comment by OVH - Friday, 25 May 2012, 02:16AM

The cut-off was "prelocalised" 800m from netcenter.
The teams are on the spot and looking for the exact location.

So for the moment we do not know the type of the cut-off and how long the reparation will take.


Comment by OVH - Friday, 25 May 2012, 02:24AM

The cut-off has been localized.Civil engineering works has cut a sleeve of 10m long. Bitumen was poured on top. The reparation is not possible.
The network teams are looking in order to find free fibers in the neighbor sleeves to make a temporary re-routing.


Comment by OVH - Friday, 25 May 2012, 02:26AM

Well ... re-routing is not possible.
The path is being re-opened to retrieve the scabbard abyss and start welding.


Comment by OVH - Friday, 25 May 2012, 04:11AM

The trench was re-opened.It's not possible to get slack on the fiber because a truck is blocking the access ...

A temporary reparation will begin. There are 144 fibers to be welded. For the moment the teams are preparing the "field".


Comment by OVH - Friday, 25 May 2012, 07:16AM

In the end, the temporary sealing has not been performed ...

The truck has been moved, access to the connexion box is available again.


Comment by OVH - Friday, 25 May 2012, 12:35PM

The weldings are in progress. Some fibers are already UP.
We still have 144 to weld.
The reestablishment deadline for everybody
is expected to 12 hours.


Comment by OVH - Friday, 25 May 2012, 12:37PM

The links have just remounted.

The NAR are pinging again.

We are waiting for the end in order to close the incident
and have more information after that.


Comment by OVH - Friday, 25 May 2012, 12:43PM


Our tube was completely weldd again and we applyed some
bitumen on it. Everything is working again.


We had an appointment yesterday with France Telecom for
an alignment problem on the securisation fiber, but
the appointment was reported to Thursday May 31st.


So the next week, we are going to set up the securisation
on the second Marseille loop. This kind of problem will
be then invisible.

We are sorry for the non availability time.